OpenStudy (anonymous):

A ball is launched from a sling shot. Its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the ball is 121 feet in the air; after 2 seconds, it is 224 feet in the air. Find the height, in feet, of the ball after 3 seconds in the air.

OpenStudy (anonymous):

@Kemist

OpenStudy (anonymous):

You say that the height is quadratic so it must be of the form h(t)=at2+bt+c. We will be able to find h at any height, and, in particular, at t= 3, if we can first find a, b, and c. To do that, we need three equations. We are told that "after 1 second, the ball is 121 feet in the air" so that h(1)= a(1)^2+ b(1)+ c= a+ b+ c= 121 and "after 2 seconds, it is 224 feet in the air" so that h(2)= a(2)^2+ b(2)+ c= 4a+ 2b+ c= 224. We need a third equation which would probably be the height at t= 0. Are you given information about that? Are we to assume that h(0)= 0 or was the slingshot initially some specifice height above the ground? Taking t= 0 we have h(0)= a(0)^2+ b(0)+ c= c. If the initial height is, in fact, 0, we have the three equations c= 0, a+ b+ c= 121, and 4a+ 2b+ c= 224. With c= 0, the last two equations become a+ b= 121 and 4a+ 2b= 224. We can eliminate b from that by subtracting twice the first equation from the second: (4a+ 2b)- (2a+ 2b)= 224- 242 or 2a= -12 or a= -6. Then a+ b= 121 becomes -6+ b= 121 so that b= 127. You now have h(t)= -6t^2+ 121t. Now, what is h(3)? If the initial height is something other than 0, just add that initial height to "h(3)".

OpenStudy (anonymous):

Not allowed to give you direct answers.